I keep getting this wrong, can someone please assist me - I will be grateful for your help!
Many graduate schools require applicants to take a certain test. Those taking the test receive three​ scores: a verbal reasoning​ score, a quantitative reasoning​ score, and an analytical writing score. In​ 2013, the average quantitative reasoning score exceeded the average verbal reasoning score by
1.1 ​points, and the average verbal reasoning score exceeded the analytical writing score by 145.7 points. The sum of the three average scores was
318.9. What was the average score for each​ category?
The average verbal reasoning score was_____
The average quantitative reasoning score was_____
The average analytical writing score was_____

Answer by addingup(3677)

analytical: x
verbal: x+145.7
quantitative: x+145.7+1.1
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
x+x+145.7+x+145.7+1.1 = 318.9
3x+292.5 = 318.9
3x = 26.4
x = 8.8 this is the analytical score
Verbal: 8.8+145.7 = 154.5
Quantitative: 8.8+145.7+1.1 = 155.6